OpenStudy (anonymous):

What is the approximate volume of a cone with a height of 6 mm and radius of 18 mm? Use 3.14 to approximate pi, and express your final answer to the nearest tenth.

OpenStudy (anonymous):

@Michele_Laino i got 2034.7 mm3 could u make sure im correct

OpenStudy (michele_laino):

you have to apply the same formula as before for previous exercise, namely: \[\Large V = \frac{{\pi \times {R^2} \times H}}{3}\] where R is the radius, and H is the height of the cone respectively

OpenStudy (anonymous):

i did

OpenStudy (michele_laino):

so we get: \[\Large V = \frac{{\pi \times {R^2} \times H}}{3} = \frac{{3.14 \times {{18}^2} \times 6}}{3} = ...?\]

OpenStudy (michele_laino):

I got: V=2034.72 mm^3

OpenStudy (michele_laino):

So your answer is correct!

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

thanks!